Sol Summary – May 5th

Crew 193 Day 13 Summary Report – 5May18

Day 13

Summary Title: It’s Been A Long Day Without You My Friend, And I’ll Tell You All About It When I See You Again (Alternate Title: Sol Long! Farewell!)

Author: Dr. Shawna Pandya, Crew Health and Safety Officer

Mission Status: Om-nom-nom-nominal

Sol Activity Summary: The day was full and busy as the Commander and XO, no longer bound by duty to travel in separate vehicles, headed toward Earth destination Grand Junction to acquire return vehicles for re-entry to civilization. Meanwhile, the intrepid crew engineer tirelessly plowed on throughout the day on a resupply mission to restock the Hab’s water’s supply. Back at the Hab, the Crew Geologist/Astronomer and Health and Safety officer (wo)manned the fort to buckle down on cleaning duties. The team reunited in the afternoon for a final scrub down of the Hab and surrounding campus, and undertook a final EVA of sorts using an ‘Earth-rover’ (also known as a car) to celebrate with a team dinner. The day can best be summarized as saying, ‘Sol’s well that end’s well…’

Look Ahead Plan:
Back to civilization in the AM!

Anomalies in work: Some unwanted visitors appeared on site, the local authorities were notified.

Weather: Partly cloudy in AM, clear and sunny in the PM.

Crew Physical Status: Nominal, if not a little crunchy on the ol’ bones here in 1g.

EVA:
A crew EVA was taken up to Marble Ritual on Rovers, for one last team bonding experience. Much merriment ensued.

Reports filed:
Sol Summary

Reports to be filed:

Operations Report

Support Requested: A gentle but loving send-off into the welcoming embrace of Earth’s dawn.

Sol Summary Report – 03 May 2018

Sol 11

Summary Title: Sol Close to the End

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Nominal

Sol Activity Summary: The entire day was spent touring around a class of 30 middle schoolers from Colorado. They viewed the Hab, the Green Hab, the Science Dome, and the Musk Observatory. They children had several good questions that our crew was able to discuss with them. They were excited about the possibilities of travelling to mars and becoming astronauts in the future! The crew enjoyed the time with the class but also long for the exploration of EVAs as they were cancelled yesterday due to rain. We intend on completing our final crew EVAs of this mission tomorrow.

Look Ahead Plan:

Tomorrow is our final day in simulation on Mars. We are looking forward to two EVAs, getting the hab ready for handover on Saturday and finishing our final reports.

Anomalies in work: None to report today, but the mud has now returned.

Weather: Partly cloudy in AM, cloudy in the PM with rain.

Crew Physical Status: Nominal

EVA:

No EVAs were completed today due to the touring class from Colorado.

Sol Summary Report – May 2nd

Crew 193 Sol Summary Report – 1MAY2018

Sol 9

Summary Title: Can’t go out cause Sol this mud.

Author: David Attig, Crew Engineer

Mission Status: Nominally muddy

Sol Activity Summary: The day started off while preparing to go out on EVA 18. Unfortunately, this was called off due to the mud outside from the rain (on Mars?) overnight. The afternoon EVA was also cancelled. The crew spent most of the day working on personal projects and relaxing. We practiced ASL some more, and enjoyed a Martian feast.

Look Ahead Plan: Tomorrow we have a group of children coming to visit us on Mars. We look forward to enlightening them on the happenings up here, and what we need to survive.

Anomalies in work:

1. Muddy mud, muddying up the muck.

Weather: Overcast and muddy in the AM, with off-and-on rain. The sun came out in the afternoon, until we were under threat of a thunderstorm (which appears to be passing us).

Crew Physical Status: Feelin Fine.

EVA: Morning EVA was cancelled today, due to inclement weather (too muddy). Afternoon EVA cancelled for same reason.

Reports filed:

GreenHab Report

Astronomy Report

EVA Request

Science Report

Sol Summary

Operations Report

EVA Report

HSO Report

Reports to be filed:

Journalist Report

Photo Report & Photo of the Day

Support Requested:

1. A Starbucks franchise on Mars.

Sol Summary – May 1st

Crew 193 Sol Summary Report – 1MAY2018

Sol 9

Summary Title: We Sol-emnly Swear We Are Up to No Good, as per the Mars-rauder’s Map

Author: Shawna Pandya, Health and Safety Officer

Mission Status: Nominally nominal

Sol Activity Summary: The day started off later than usual as the EVA-16 team opted to delay their morning excursion until 0900 in the face of threatening storm clouds to the South and East. The morning EVA then progressed as expected, with the PM team keeping up with usual Hab operations, cleaning schedules, GreenHab upkeep and data collection in the interim. Lunch was a Martian feast of curried kidney beans and brown rice (so much better than chianti and fava beans). The EVA-17 team left shortly thereafter to scout the area just west of Zubrin’s head, wisely choosing to turn back 90min ahead of schedule in anticipation of the weather, luckily only meeting with Martian rain for a slight period at the tail end of their sojourn. The crew astronomer managed to capture one solar prominence and a filament. Hab operations otherwise continued as per usual, with one brave crew member opting to try waterless shampoo, and photo-documenting the results (spoiler alert: said crew member still has all her hair).

Look Ahead Plan: EVA 18 will consist of a Green Valley and Mid-Ridge Planitia exploration, whereas EVA 19 will take up an excursion just Southwest of Kissing Camels Ridge. The team continues to function well under the guidance of the dear Commander. Crew development and team-building will continue by way of ASL, team stories, movies and a good ol’ fashioned round of Shanghai, although to be fair, the latter is probably more likely to result in team-breaking than anything else.

Anomalies in work:

1. The crew astronomer was accidentally locked out for a brief spell, but was let in after pleading his case by way of loud noises, and harbors no grudges (we hope).

Weather: Overcast in the AM, rainy and gusty (a Mars-nado, really) in the PM, resulting in a premature return to base on PM EVA.

Crew Physical Status: "So good it hurts," as per one team member.

EVA: EVA-16 proceeded with an exploration of Lith Canyon to the North, and accrued the usual data regarding emergency shelters and crew comfort, safety, health and well-being. Similarly, EVA-17 proceeded with an exploration just West of Zubrin’s Head, and accrued the same data regarding shelters and crew well-being. EVA-17 was halted prematurely due to incrementally inclement weather, which then became overtly inclement.

Reports filed:

GreenHab Report

Astronomy Report

EVA Request

Journalist Report

Science Report

Sol Summary

Photo Report & Photo of the Day

Operations Report

Reports to be filed:

EVA Report

Support Requested:

1. Yogurt. Please send all the yogurt. The crew’s blood levels of Vitamin Y are critically low, and it makes us sad. 🙁

Sol 8 Summary – April 30th

Crew 193 Sol Summary Report – 30 APR 2018

Sol 8

Summary Title: “It was a good Sol” – Ice Cube

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Nominal

Sol Activity Summary: The day started off later than normal as the first EVA was not scheduled to leave until 0800. EVA 14 was able to explore the White Moon area on Mars! While EVA 14 was occurring, the two crew members in the hab took care of the plants in the green Hab and cleaned the shared living area in the main Hab. One crew member credits the growth of the plants to the time she spends talking to them. Maybe Mark Watney should have tried that!

The crew lunched together (thanks again to our commander for cooking), had a short meeting to continue team development and then it was time to get ready for EVA 15. EVA 15 was cut short at 1800 due to approaching rain clouds. The EVA was in the Candor Chasma area and it was decided that if it rained it would be a very difficult area to exit and return to the Hab.

Look Ahead Plan: EVA 16 will be exploring Lith Canyon and the Dinosaur Quarry. EVA 17 will explore the area around Zubrin’s Head and Southwest. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better. Our daily team meetings have been working well and will be continued through the remainder of the mission.

Anomalies in work:

1. None to report today

Weather: Partly cloudy in AM, cloudy and gusty in PM with rain in the distance.

Crew Physical Status: Minor physical ailments (blister, etc.) but otherwise nominal!

EVA:

EVA 14 – Explored the moon area. Continued the GPS study and improved on their ATV skills by taking their most challenging road off of Cowdung Road yet.

EVA 15 – Successfully explored the Candor Chasma area and obtained some spectacular photographs. EVA was cut short early due to the possibility of incoming rain making it difficult to exit the Candor Chasma area.

**Please note that the EVA numbers submitted in the EVA report are not correct. It should have been EVA 14 and 15, not 15 and 16**

Reports filed:

GreenHab

Science Report

Astronomy Report

Operations Report

EVA Report

Reports to be filed:

Commander’s Report

Photo & Photo of the Day Report

Next Day EVA Request

Support Requested:

1. Hot Tub? (No requests at this time)

Sol Summary – April 29th

Crew 193 Sol Summary Report – 29APR2018

Sol 7

Summary Title: How Windy Is It? It’s Sooooollllll Windy, the Answer Will Blow You Away!

Author: Dr. Shawna Pandya, HSO

Mission Status: Supra-nominally windy.

Sol Activity Summary:
Today the two EVA teams swapped AM/PM spots to change things up, and will be doing so for the rest of their time here. EVA-AM team explored White Rock Canyon and Kissing Camels Ridge, and a good time was had by all, other than a sprained knee sustained by one of the team, who is now recovering. Meanwhile, EVA-PM team took care of the GreenHab and wo/manned the fort. Upon her return, the Commander took Command of the kitchen and whipped up a delicious feast of pasta alfredo with chicken, spaghetti bolognaise and a hearty vegetable soup to whet her picky crew’s carnivorous, non-dairy and vegetarian predilections, respectively. In the late afternoon, EVA Team 2 took to the mean streets of Mars for a sojourn up North, certain that they would have turn back early due to mounting winds and increasing cloud cover. Alas, the winds of Mars made way for our fearless explorers, who had the time (and view) of their lives up by Copernicus way. In the midst of these EVAs, more scientific data was accrued and solar prominences observed by the science team and astronomer. The Sol will culminate with another hearty Martian spread come dinnertime, followed by some ‘Gold’-en opportunities for crew bonding, by way of ASL, cards and games.

Look Ahead Plan:
The week continues with more EVAs to the Moon by the EVA-AM team, and a PM trek over to Cactus Road with exploration of the near-by sights and sounds, with plans to forge on with the usual science, Hab operations and astronomy.

Anomalies in work:

1. Generator is not charging solar panels overnight – ongoing issue, no changes.
2. Suit #10 has been decommissioned due to aberrant battery contacts.

Weather: Partly overcast, very windy, warm.

Crew Physical Status: One crew member sprained a ligament on EVA and is recovering. They have plans to go on a light excursion EVA (light walking on flat terrain or mostly in vehicle), or plans to stay back if not recovered in the morning.

EVA:

Reports filed:
EVA Report
Astronomy Report

Science Report
Operations Report

HSO Report
GreenHab Report
Sol Summary

Reports to be filed:
GreenHab
Photo & Photo of the Day Report

Commander’s Report
EVA Request

Support Requested:
1. Straps to 2-piece suit neck ring
2. New terminal connectors for Suit #10

Sol 6 Summary Report – April 28th

Sol 6

Summary Title: Stopping to Smell the Roses

Author: Eric Shear, CGA

Mission Status: Perfectly nominal… for now.

Sol Activity Summary: This morning we did a series of short outreach videos per the astronomer’s suggestion. For lunch, we feasted on (yet) another delicious Martian lunch of quinoa and curry chicken courtesy of our commander’s mean cooking skills. The afternoon was spent on ongoing science experiments, finishing the MDRS sign, building the new red wagon for the Greenhouse, and usual Hab operations. Our crew astronomer managed to catch a small prominence on the sun’s limb during his observations this morning. The remainder of the day was spent planning out our remaining EVAs. We now have a week’s worth of EVA destinations for both away teams.

Look Ahead Plan: Tomorrow, we will execute the first two EVAs of the new week. In the morning, the American team will head to Zubrin’s Rock and explore White Rock Canyon. Team Canada will follow up at Toothy Ridge in the afternoon.

Anomalies in work:

None

Weather: Cloudy skies in AM and PM. Extremely hot.

Crew Physical Status: Nominal

EVA: None

Reports filed:

Astronomy Report

GreenHab Report

Reports to be filed:

EVA Report

EVA Request

Science Report

Journalist Report

Photo & Photo of the Day Report

Operations Report

Support Requested:

None

Sol Summary – April 27th

Crew 193 Sol Summary Report – 27APR2018

Sol 5

Summary Title: All Aboard the Sooooollllll-Train

Author: Dr. Shawna Pandya, HSO

Mission Status: Nominal like you wouldn’t believe. Bigly Nominal. My friends tell me they’ve never seen a mission more nominal.

Sol Activity Summary: Today we planned for a shorter set of EVAs in anticipation of the hot day and gusting winds, concentrating on medical skills and simulation (see below). For lunch, we feasted on a delicious Martian lunch of Jamabalaya rice, broccoli and cheese soup, and fish. The afternoon was then spent on testing new radios and equipment, ongoing science experiments, finishing the MDRS sign, and usual Hab operations. Notably, our crew astronomer did a stellar job of picking up a solar flare, leading to the declaration of a state of emergency from the ensuing radiation storm, and reinforcing the good work we are doing scouting and mapping potential emergency shelters while on EVA. The remainder of the day was spend planning out our remaining week on Mars.

Look Ahead Plan: Tomorrow will be a mandatory rest day with no EVAs planned in anticipation of the 30C high. We will plan out the rest of our time here at MDRS, including a school visit on Sol 11, May the 4th on Sol 12, and end-of-rotation operations on Sol 13.

Anomalies in work:

1. Paintbrushes were not small enough for lettering on MDRS sign.
2. Soldering iron is broken.

Weather: Clear skies in AM, extremely hot by mid-day.

Crew Physical Status: Affected crew member from EVA recovering nicely.

EVA: The crew spent the morning going over principles of wilderness medical triage, evacuation and critical decision-making in case of accident while on EVA. We then moved these skills outside in our spacesuits, simulating various medical emergencies, including extraction from a rolled ATV while maintaining C-spine precautions (easier said than done in a spacesuit!). Our second EVA was cut short due to heat exhaustion in one crew member.

Reports filed:
EVA Report
EVA Request
Astronomy Report

Science Report
Journalist’s Report

Reports to be filed:
GreenHab
Photo & Photo of the Day Report

Operations Report

Support Requested:
1. Smaller paint brushes for MDRS signs.

Sol Summary Report – 26 April 2018

Sol 4

Summary Title: Broken Glasses and Martian Pizza

Author: David Attig, Crew Engineer

Mission Status: Super Tremendous Nominal.

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 8 which left at 0730.

The crew feasted on some homemade pizza, with some fresh salad and soup made with our greenhouse produce. We spent time on personal work and research projects before EVA 9.

While EVA 9 was underway, the HAB was vacuumed and progress made on experiments.

Look Ahead Plan: Tomorrow we will stick close to the Hab due to predicted hot weather and high winds in the afternoon. We will have two EVAs to practice on ATVs, and will spend the afternoon in the Hab practicing medical emergency situations and reviewing fire safety.

Anomalies in work:

1. There was a pair of sunglasses lost during the afternoon EVA. They were found, but had been driven over by a Rover.

2. There was a missed picture opportunity; the crew member who was asked to take it did a poor job and the crew member who asked for it did not review the picture before returning from EVA.

Weather: Clear skies in AM, chilly in morning and quickly warming during the day. Warm and breezy in PM.

Crew Physical Status: Nominal

EVA:

EVA 8 – went well, gray moon area was explored. There is incredible scenery, the landscape matches up with its name very well. Upon returning, they brought materials in from the RAM to work on in the HAB.

EVA 9 – Saw a fantastic view of Candor Chasma.

Sol Summary Report April 25th

Crew 193 Sol Summary Report – 25 APR 2018

Sol 3

Summary Title: Life Will Find a Way

Author: Doug Campbell, Executive Officer and Science Officer

Mission Status: Sub-optimal due to minor crew health issues but expected to be back to nominal tomorrow

Sol Activity Summary: The day started off early with quick individual breakfasts in order to prepare for EVA 6 which left at 0730. EVA 6 was a successful failure in that valuable exploration of Lith Canyon was completed; however, the intended destination was the White Moon area! While EVA 6 was occurring, the plants were watered in the Green Hab and some vegetables were harvested. Our crew astronomer made a quick trip to the observatory and did some troubleshooting with the GPS units. Our crew engineer did work around the hab and communicated with mission control.

The crew lunched together (thanks to our commander for cooking!) and then spent time on personal work and research projects before EVA 7. EVA seven was cut short at 1600 by the Health and Safety Officer over concerns of crew dehydration. Late afternoon was spent on reports. The plan for tonight is to watch a movie as a team, perhaps…. The Martian!!

Look Ahead Plan: EVA 8 will attempt to complete the mission plan that was submitted for EVA 6 but not completed. EVA 9 will be exploring further into Candor Chasma as part of the GPS mapping study. We will continue to work on learning ASL for 5 minutes each day as well as sharing stories about our lives so that the crew gets to know each other better.

Anomalies in work:

1. Radio #6 has a poor battery which loses charge quickly. It caused Gold 001 to be without communication during EVA 6. This radio has been put out of service by crew 193
2. EVA 7 was cut short due to health concerns over crew dehydration

Weather: Clear skies in AM, partly cloudy and gusty in PM.

Crew Physical Status: Some blisters, one incidence of crew mild dehydration. All members are resting and back to nominal by days end.

EVA:
EVA 6 – Successfully explored the edges and the base of Lith Canyon. As per the title of this report, the EVA team found group of fully blossomed wild flowers on the edge of a large, sandy rock formation with no other flora or fauna in sight; proving that even on mars, life will find a way.
EVA 7 – Began the EVA in the RAM to paint the new sign for the MDRS campus. The crew then proceeded to climb the hill up to hab ridge. However, during the climb, one of the members began to not feel well and the remainder of the EVA was scrubbed.

Reports filed:
GreenHab
Science Report
Astronomy Report
Journalist’s Report
Photo & Photo of the Day Report

Reports to be filed:
Operations Report
EVA Report

Support Requested:

Additional fuel for ATVs.
Orange or green paint for the MDRS sign (or both)
Hot pads so we can put hot cookware on the table without wrecking the tabletop.